OpenStudy (anonymous):

Please Help! Medal! Solve the proportion. Where necessary, round to the nearest hundredth. 8/2 = x+3/4 A. 13 B. 1 C. 7 D. 16 My answer: D

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

x = 13/4

OpenStudy (anonymous):

i think A

OpenStudy (appleduardo):

(8/2) - (3/4) = x 13/4 = x :p

OpenStudy (bibby):

set up the proportion right. \(\huge \frac{8}{2}=\frac{ x+3}{4}\) 2(x+3)=32 2x+6=32